in what ratio does the line 2x-3y=2 divide the line segment join of (4,6) and (8,1)?

Grade:9

1 Answers

Shreya shetty
5 years ago
Given :a=2,b=-3,c=-2,and (x1,y1)=(4,6) (x2,y2)=(8,1)
The given line ax+by+c=0 devides the line joining the points (x1,y1) and (x2,y2) in the ratio =-(ax1+by1+c)/(ax2+by2+c)
Therefore, -(2(4)+(-3)(6)+(-2))/(2(8)+(-3)1+(-2))=-(-12)/11= 12/11